Samuel has found that his new car gets 37 miles per gallon on the highway and 30 miles per gallon in the city. He recently drove 365 miles on 11 gallons of gasoline. How many miles did he drive on the highway? How many miles did he drive in the city
Answer by CubeyThePenguin(3113)

H = gallons used driven on highway
C = gallons used driven in city

37H + 30C = 365
H + C = 11

(H, C) = (5, 6)

Samuel drove 6 * 30 = 180 miles in the city.